# Edge Modes on Stringy Horizons

### v1

@@ I Introduction
A Euclidean path integral $\hat{Z}(\beta)$ on a manifold $S^{1}\times M$ admits a natural Lorentzian interpretation as a thermal partition function. One can regard $M$ as the spatial slice and the generator of time translations along $S^{1}$ as the canonical Hamiltonian $H$. The path integral can then be identified with the partition function $\mbox{Tr}_{\mathcal{H}}[e^{-\beta H}]$, where the inverse temperature $\beta$ is the length of the $S^{1}$ and $\mathcal{H}$ is the Hilbert space of wave-functions defined over $M$. A path integral over a generic Euclidean manifold does not admit such a canonical thermal interpretation.
@@
The physical idea behind this split is the following. One could try to identify the generator of time translations in the static dS patch as the canonical Hamiltonian. Observers in this patch experience a heat bath at inverse temperature $2\pi$, suggesting a thermal interpretation. However, this generator on the Euclidean manifold has a fixed point corresponding to the bifurcate horizon in the Lorentzian continuation. As a result, the spatial slicing is not smooth at the fixed point, leading to a correction term from the modes localized near the horizon.
@@
$\log\hat{Z}_{\text{edge}}:=\sum_{i}\log\hat{Z}^{i}_{\text{edge}}=Z_{\text{edge}}\,$, (2)
where we have identified the logarithm of the spacetime partition function with the worldsheet partition function $Z_{\text{edge}}$ for the edge modes, extending the usual identification for the bulk modes. We use the ‘hat’ for spacetime partition functions to distinguish from the worldsheet partition functions with no hat. Using the heat kernel representation of the spacetime determinants and identifying the Schwinger parameters with modular parameters, one can obtain a representation of the worldsheet partition function as a modular integral. This is the object that we analyze and show that it naturally leads to a modular invariant expression.
@@ II Edge modes in Quantum Field Theory
where $H$ is the generator of global time translations, and $\mbox{tr}_{G}$ is the trace over the global single-particle Hilbert space that furnishes the representation. The equivalence between $\mbox{tr}_{S}$ and $\mbox{tr}_{G}$ is a consequence of a one-to-one mapping between the single-particle states with the same $\omega$ in the static and global patches via the Bogoliubov transformation. For a massive integer spin-$s\geq 1$ field, the bulk Harish-Chandra character is given by
@@
where $D^{(d)}_{s}$ is the $SO(d)$ spin degeneracy.
Interestingly, the edge contribution also turns out to be given in terms of a Harish-Chandra character but in two fewer dimensions [2]:
$\chi^{\text{dS}}_{\text{edge}}(u)=D^{(4)}_{s-1}\,\left(e^{-iu\nu}+e^{iu\nu}\right)\,.$ (7)
As shown in [2], integrating over $\chi^{\text{dS}}_{\text{bulk}}(u)$ as in (4) gives $\log\hat{Z}^{i}_{\text{bulk}}$ which equals the logarithm of the thermal partition function. Thus, it is clear from (4) that for a generic bosonic field with spin, the Euclidean path integral is not exactly equal to the thermal partition function. A correction term is needed, which can be identified with the contribution from the edge modes, $\log\hat{Z}_{\text{edge}}$ in (1).
The representation above in terms of the characters is useful to see the split between the bulk and the edge modes from the Lorentzian perspective. This split can also be seen from the Euclidean perspective [2], by noting
@@
using $D^{(2)}_{s}=D^{(2)}_{n+1}=1$, since all $SO(2)$ representations are one-dimensional. The extra factor of $2$ for the first term arises from the fact that a spin $s$ field contains the contributions from both helicities $\pm s$. Substituting into (3), the $2(n+1)^{2}$ term gives rise to the bulk character, and the $s^{2}$ piece to the edge character, which equals the character of $s^{2}$ scalar fields in two lower dimensions. This group theoretic formulation has a physical interpretation in terms of quasi-normal modes [2, 39], which could be useful in less symmetric backgrounds [33, 27].
@@ III Edge Modes in String Theory
$Z_{\text{bulk}}=\log\hat{Z}_{\text{bulk}}=\sum_{i}\log\hat{Z}_{\text{bulk}}^{i}\,.$ (11)
@@
$Z_{\text{bulk}}=\frac{L^{3}R^{23}}{\ell^{26}}\int_{\mathcal{S}}\frac{d^{2}\tau}{\tau_{2}^{13}}\mbox{tr}{}\,\left[q^{N-1}\bar{q}^{\tilde{N}-1}\right]\,$, (13)
@@
$\begin{split}Z_{\text{edge}}\sim-\frac{L\,R^{23}}{\ell^{24}}\int_{0}^{\infty}\frac{d\tau_{2}}{\tau_{2}^{12}}\sum_{i}\,s_{i}^{2}\,e^{-\pi\ell^{2}\tau_{2}m_{i}^{2}}\,.\end{split}$ (17)
In the flat limit, we have three real worldsheet bosons with $\mathbb{R}^{3}$ as the target space. We combine two of them into a complex boson $X:=X_{1}\pm iX_{2}$ with unit charge under the $U(1)$ generator corresponding to the rotation in the plane. The remaining direction $X_{3}$ is uncharged. The quantum number $s$ corresponds to the spin in the 1-2 plane, excluding the orbital angular momentum coming from the zero modes. Summing over species in (17) implies inserting an operator $\frac{1}{2}(S+\tilde{S})^{2}$ in the worldsheet trace, where $S$ and $\tilde{S}$ are the spin operators of the left and right movers. The factor of half accounts for the fact we sum over both $\pm s$ helicities.
@@
$Z_{\text{edge}}\sim-\frac{L\,R^{23}}{\ell^{24}}\int_{0}^{\infty}\frac{d^{2}\tau}{\tau_{2}^{12}}\mbox{tr}{}\,\left[(S+\tilde{S})^{2}\,q^{N-1}\bar{q}^{\tilde{N}-1}\right]\,.$ (18)
@@
It implies that $E(\tau,s)\sim sE(\tau,1-s)$ near $s=0$. Thus, instead of $s=0^{+}$ we can examine the behavior near $s=1^{-}$. Precisely, this Laurent expansion is given by the classic Kronecker limit formula [32]:
@@ IV Physical Interpretation
Edge modes arise in gauge theory due to gauge constraints such as Gauss’ law for a Maxwell field [24, 25, 21, 22, 5]. For a Proca field in a spontaneously broken gauge theory, the necessity for edge modes can be seen most clearly by quantizing the Fierz-Pauli system [7, 2]. String theory has a large (possibly spontaneously broken) gauge invariance, with an infinite tower of massive fields. The partition function (38) captures the edge modes for this tower. For each mass $m$ and helicity $r$, one obtains a unitary irreducible representation $[m,r]$ of the Poincaré group $ISO(1,2)$. CPT invariance requires that, for a given spin $s$, both helicities $r=\pm s$ must belong to the spectrum. The representation $m_{i},s_{i}]\oplus[m_{i},-s_{i}]$ can be obtained as a Hilbert space by quantizing a Fierz-Pauli theory with Stückelberg fields for a totally symmetric tensor $\Phi_{\mu_{1}\mu_{2}\ldots\mu_{s}}$ which is transverse and traceless ($\nabla^{\nu}\Phi_{\nu\mu_{1}\ldots\mu_{s-1}}=\Phi^{\nu}_{\nu\mu_{1}\ldots\mu_{s-2}}=0$), described by a local gauge-invariant action.
@@
Instead of computing the edge contributions near the Rindler horizon directly, we have considered its deformation to the de Sitter horizon with extremely low curvature. The states now furnish representations $\nu_{i},s_{i}]\oplus[\nu_{i},-s_{i}]$ in the principal series of the de Sitter group $SO(1,3)$, regarding the Poincaré group as a Wigner-Inönü contraction of the de Sitter group. Using the Faddeev-Popov procedure for the path integral, one obtains the edge mode contribution from the ‘unmatched’ zero modes of ghost fields that do not fully cancel against the contribution from the Stückelberg fields, consistent with locality and unitarity [7, 2].
@@
A short proper time expansion of the character integral (4) is in agreement with the known heat kernel coefficients [41]. The edge corrections could be interpreted as ‘contact terms’, as suggested by the heuristic picture of Susskind and Uglum [40] of a string worldsheet at the horizon and computed for the photon and the graviton field in [31, 6]. The contact terms arise from couplings linear in the Riemann tensor in the quadratic fluctuation operator and are generically present for any spin-$s\geq 1$ field. We emphasize that the modular invariant expression (38) is free of UV divergences. Moreover, the integral over the proper time contains a lot more information than the short proper-time expansion, including the nonlocal and finite contributions from the infinite tower of Kaluza-Klein modes in the large $R$ limit.
@@
The edge modes contribute to the entanglement entropy associated with the horizon as the entangling surface. They could also be related to soft modes [3]. The formal expression (4) for $Z_{\text{edge}}$ in quantum field theory is UV divergent. As we have seen, after summing over the string tower, one obtains a UV convergent expression (38) because of modular invariance. For the bosonic string there is the usual tachyonic IR divergence, which is expected to be absent for the superstring and will be explored elsewhere. One expects that the total entanglement entropy will receive contributions also from the bulk term, corresponding to the area divergence due to the divergent local temperature of the thermal bath near the horizon. These questions will be explored in [12]. Independently of its relation to entanglement entropy, $Z_{\text{edge}}$ is interesting in its own right, since it is related to the underlying gauge symmetry.
@@
It is useful to formulate these results in the language of algebraic quantum field theory. In this formalism there is no UV cut-off. The UV divergence of $\hat{Z}_{\text{edge}}$ (or of entanglement entropy) corresponds to the fact that, unlike in a bipartite quantum mechanical system, the algebra of observables in a QFT is Type-III rather than Type-I. See [42] for a review. With the inclusion of gravity, the algebra of observables can become Type-II [44, 10], a fact possibly related to the renormalization of Newton’s constant in canonical gravity [40]. However, one would really like to show that the algebra is more like Type-I. This is expected also from the holographic description of the two-sided black hole, since the algebra of observables acting on a single copy of the dual CFT furnishes an irreducible representation on the Hilbert space of the CFT and is thus Type-I. It is not clear how to define local observables in string theory but the UV finiteness of $\hat{Z}_{\text{edge}}$ is a strong indication of a Type-I behavior.

### v2

@@ I Introduction
A Euclidean path integral on a manifold $S^{1}\times M$ admits a natural Lorentzian interpretation as a thermal partition function. One can regard $M$ as the spatial slice and the generator of time translations along $S^{1}$ as the canonical Hamiltonian $H$. If $\beta$ is the length of the $S^{1}$, then the path integral $\hat{Z}(\beta)$ can be identified with the partition function $\mbox{Tr}_{\mathcal{H}}[e^{-\beta H}]$ where $\mathcal{H}$ is the Hilbert space of wave-functions defined over $M$. A path integral over a generic Euclidean manifold does not admit such a canonical thermal interpretation.
@@
The physical origin of this split is the following. One could try to identify the generator of time translations in the static dS patch as the canonical Hamiltonian. Observers in this patch experience a heat bath at inverse temperature $2\pi$, suggesting a thermal interpretation. However, this generator on the Euclidean manifold has a fixed point corresponding to the bifurcate horizon in the Lorentzian continuation. As a result, the spatial slicing is not smooth at the fixed point, leading to a correction term from the modes localized near the horizon.
@@
$Z_{\text{edge}}:=\log\hat{Z}_{\text{edge}}:=\sum_{i}\log\hat{Z}^{i}_{\text{edge}}\,$, (2)
where we have identified the worldsheet partition function $Z_{\text{edge}}$ with the logarithm of the total spacetime partition function for the edge modes, obtained by summing over all strings fields, extending the usual identification for the bulk modes. We use the ‘hat’ for spacetime partition functions to distinguish from the worldsheet partition functions with no hat. Using the heat kernel representation of the spacetime determinants and identifying the Schwinger parameters with modular parameters, one can obtain a representation of the worldsheet partition function as a modular integral. This is the object that we analyze and show that it naturally leads to a modular invariant expression.
@@ II Edge modes in Quantum Field Theory
where $H$ is the generator of global time translations in the static patch, and $\mbox{tr}_{G}$ is the trace over the global single-particle Hilbert space that furnishes the representation. The equivalence between the trace over single-particle states in the static patch and $\mbox{tr}_{G}$ is a consequence of a one-to-one mapping between states with the same $\omega$ in the static and global patches via the Bogoliubov transformation. For a massive integer spin-$s\geq 1$ field, the bulk Harish-Chandra character is given by
@@
where $D^{(d)}_{s}$ is the $SO(d)$ spin degeneracy. In particular, $D^{(2)}_{s}=2$, since all $SO(2)$ representations are one-dimensional and a spin $s$ field contains the contributions from both helicities $\pm s$. Note also that $D^{(4)}_{s}=(s+1)^{2}$.
Interestingly, the edge contribution also turns out to be given in terms of a Harish-Chandra character, but in two fewer dimensions [2]:
$\chi^{\text{dS}}_{\text{edge}}(u)=D^{(4)}_{s-1}\,\left(e^{-iu\nu}+e^{iu\nu}\right)\,$, (7)
As shown in [2], integrating over $\chi^{\text{dS}}_{\text{bulk}}(u)$ as in (4) gives $\log\hat{Z}^{i}_{\text{bulk}}$ which equals the logarithm of the thermal partition function. Thus, it is clear from (4) that for a generic bosonic field with spin, the Euclidean path integral is not exactly equal to the thermal partition function. A correction term is needed, which can be identified with the contribution from the edge modes, $\log\hat{Z}^{i}_{\text{edge}}$ in (1).
The representation above in terms of the characters is useful to see the split between the bulk and the edge modes from the Lorentzian perspective. This split can also be seen from the Euclidean perspective [2], by noting that for $n\geq 0$,
@@
while $D^{(4)}_{-1,s}=-s^{2}$. Substituting into (3), the $2(n+1)^{2}$ term gives rise to the bulk character, and the $s^{2}$ terms to the edge character, which equals the character of $s^{2}$ scalar fields in two lower dimensions. This group theoretic formulation has a physical interpretation in terms of quasi-normal modes [2, 40], which could be useful in less symmetric backgrounds [34, 28].
@@ III Edge Modes in String Theory
$Z_{\text{bulk}}:=\log\hat{Z}_{\text{bulk}}:=\sum_{i}\log\hat{Z}_{\text{bulk}}^{i}\,.$ (11)
@@
$Z_{\text{bulk}}=\frac{L^{3}R^{23}}{\ell^{26}}\int_{\mathcal{S}}\frac{d^{2}\tau}{\tau_{2}^{14}}\mbox{tr}{}\,\left[q^{N-1}\bar{q}^{\tilde{N}-1}\right]\,$, (13)
@@
$\begin{split}Z_{\text{edge}}\sim-\frac{L\,R^{23}}{\ell^{24}}\int_{0}^{\infty}\frac{d\tau_{2}}{\tau_{2}^{13}}\sum_{i}\,s_{i}^{2}\,e^{-\pi\ell^{2}\tau_{2}m_{i}^{2}}\,.\end{split}$ (17)
In the flat limit, we have three real worldsheet bosons with $\mathbb{R}^{3}$ as the target space. We combine two of them into a complex boson $X:=X_{1}\pm iX_{2}$ with unit charge under the $U(1)$ generator, corresponding to the rotation in the plane. The remaining direction $X_{3}$ is uncharged. The quantum number $s$ corresponds to the spin in the 1-2 plane, excluding the orbital angular momentum coming from the zero modes. Summing over species in (17) implies inserting an operator $\frac{1}{2}(S+\tilde{S})^{2}$ in the worldsheet trace, where $S$ and $\tilde{S}$ are the spin operators of the left and right movers. The factor of half accounts for the fact we sum over both $\pm s$ helicities.
@@
$Z_{\text{edge}}\sim-\frac{L\,R^{23}}{\ell^{24}}\int_{0}^{\infty}\frac{d^{2}\tau}{\tau_{2}^{13}}\mbox{tr}{}\,\left[(S+\tilde{S})^{2}\,q^{N-1}\bar{q}^{\tilde{N}-1}\right]\,.$ (18)
@@
It implies that $E(\tau,s)\sim sE(\tau,1-s)$ near $s=0$. Thus, instead of $s=0^{+}$ we can examine the behavior near $s=1^{-}$. Precisely this Laurent expansion is given by the classic Kronecker limit formula [33]:
@@ IV Physical Interpretation
Edge modes arise in gauge theory due to gauge constraints, such as Gauss’ law for a Maxwell field [25, 26, 22, 23, 5]. For a Proca field in a spontaneously broken gauge theory, the necessity for edge modes can be seen most clearly by quantizing the Fierz-Pauli system [7, 2]. String theory has a large (possibly spontaneously broken) gauge invariance, with an infinite tower of massive fields. The partition function (38) captures the edge modes for this tower. For each mass $m$ and helicity $r$, one obtains a unitary irreducible representation $[m,r]$ of the Poincaré group $ISO(1,2)$. CPT invariance requires that, for a given spin $s$, both helicities $r=\pm s$ must belong to the spectrum. The representation $m_{i},s_{i}]\oplus[m_{i},-s_{i}]$ can be obtained as a Hilbert space by quantizing a Fierz-Pauli theory with Stückelberg fields for a totally symmetric tensor $\Phi_{\mu_{1}\mu_{2}\ldots\mu_{s}}$ which is transverse and traceless ($\nabla^{\nu}\Phi_{\nu\mu_{1}\ldots\mu_{s-1}}=\Phi^{\nu}_{\nu\mu_{1}\ldots\mu_{s-2}}=0$), described by a local gauge-invariant action.
@@
Instead of computing the edge contributions near the Rindler horizon directly, we have considered its deformation to the de Sitter horizon with extremely low curvature. The states now furnish representations $\nu_{i},s_{i}]\oplus[\nu_{i},-s_{i}]$ in the principal series of the de Sitter group $SO(1,3)$, regarding the Poincaré group as a Wigner-Inönü contraction of the de Sitter group. Properly taking into account the contribution from the ‘unmatched’ zero modes of the Faddeev-Popov ghost fields, the bulk and edge contributions organize into Harish-Chandra characters (4) consistent with locality and unitarity [7, 2].
@@
A short proper time expansion of the character integral (4) is in agreement with the known heat kernel coefficients [42]. The edge corrections could be interpreted as ‘contact terms’, as suggested by the heuristic picture of Susskind and Uglum [41] of a string worldsheet at the horizon and computed for the photon and the graviton field in [32, 21, 6]. The contact terms arise from couplings linear in the Riemann tensor in the quadratic fluctuation operator and are generically present for any spin-$s\geq 1$ field. We emphasize that the modular invariant expression (38) is free of UV divergences. Moreover, the integral over the proper time contains a lot more information than the short proper-time expansion, including the nonlocal and finite contributions from the infinite tower of Kaluza-Klein modes in the large $R$ limit.
@@
The edge modes contribute to the entanglement entropy associated with the horizon as the entangling surface. They could also be related to soft modes [3]. The formal expression (4) for $\hat{Z}_{\text{edge}}$ in quantum field theory is UV divergent. As we have seen, after summing over the string tower, one obtains a UV convergent expression (38) because of modular invariance. For the bosonic string there is the usual tachyonic IR divergence, which is expected to be absent for the superstring. Moreover, one expects that the total entanglement entropy will receive contributions also from the bulk term, corresponding to the area divergence due to the divergent local temperature of the thermal bath near the horizon. These questions will be explored in [12]. Independently of its relation to entanglement entropy, $Z_{\text{edge}}$ is interesting in its own right, since it is related to the underlying gauge symmetry.
@@
It is useful to formulate these results in the language of algebraic quantum field theory. In this formalism there is no UV cut-off. The UV divergence of $\hat{Z}_{\text{edge}}$ (or of entanglement entropy) corresponds to the fact that, unlike in a bipartite quantum mechanical system, the algebra of observables in a QFT is Type-III rather than Type-I. See [43] for a review. With the inclusion of gravity, the algebra of observables can become Type-II [45, 10], a fact possibly related to the renormalization of Newton’s constant in canonical gravity [41]. However, one would really like to show that the algebra is more like Type-I. This is expected also from the holographic description of the two-sided black hole, since the algebra of observables acting on a single copy of the dual CFT furnishes an irreducible representation on the Hilbert space of the CFT, and is thus Type-I. It is not clear how to define local observables in string theory but the UV finiteness of $Z_{\text{edge}}$ is a strong indication of a Type-I behavior.